A 700.0 mL solution with 0.20 moles of ammonia (Kb = 1.8x10-5) is made. Solve for the pH of each solution below after it is added to our initial ammonia solution.

(a) 700.0 mL of pure water.

(b) 700.0 mL of pure water and 0.20 moles of hydrochloric acid.

(c) 700.0 mL of pure water and 0.30 moles of hydrochloric acid.
Answer #1

NH₂-weak base K6 21-88105 pks --log K6 = -log €48 x105) 24.74 ® 700 me pure wated Total volume = 200+700= 1. = 104 lit [NH3]